40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 1 00:05:03 2013, Koji, Update, LSC, More stable lock of PRMI (REF33I and AS55Q) Screenshot.png
    Reply  Wed May 1 10:42:35 2013, Steve, Update, LSC, cleanup  
    Reply  Wed May 1 14:42:45 2013, Jenne, Update, LSC, POP now has lens in front of PD 
Message ID: 8519     Entry time: Wed May 1 14:42:45 2013     In reply to: 8517
Author: Jenne 
Type: Update 
Category: LSC 
Subject: POP now has lens in front of PD 

Quote:

- At the end of the session, Jenne told me that the POP PD still has a large diameter beam. (and a steering mirror with a peculiar reflection angle.)
==> THIS SHOULD BE FIXED ASAP
because the normalization factor can be too much susceptible to the misalignment of the spot.

 Koji set the IFO in a PRM-ITMY configuration for me, while I went to put a lens on the POP path.  Before putting the lens, the maximum average output that I saw from the diode (on a 'scope) was 4.40mV.  After putting in the lens and realigning the beam onto the diode, the new max DCvalue that I saw was 21.6mV.  This is a factor of 4.9. 

EDIT:  The dark value was -3.20mV, so actually the ratio is ~3.25 .

I have not yet done anything to fix the situation of the large angle of incidence on the first out-of-vac steering mirror.

ELOG V3.1.3-